# Robust Preference Optimization through Reward Model Distillation

### v1

@@ 5 Theoretical analysis
The optimum in Proposition 3 corresponds to $\log\psi_{w}/\psi_{\ell}={\beta}^{-1}\log(\alpha-1)$. Recall that IPO seeks to assign a constant value to this ratio by minimizing $(\log\frac{\psi_{w}}{\psi_{\ell}}-\tau^{-1})^{2}$; the (unconstrained) optima are identical for $\tau^{-1}:={\beta}^{-1}\log(\alpha-1)$, but the loss surfaces are different (see Appendix B). DPO sets $\pi_{\theta}(y^{\ell}_{i})\to 0$, as shown in Corollary 1; this is due not only to competition from $\pi_{\theta}(y^{w}_{i})$ but from DPO penalizing positive probability on $y^{\ell}_{i}$. Analysis of the distilled loss gives a similar result:
@@
For any fixed $\hat{\pi}_{\theta}(y_{i}^{w})$ and $\beta>0,$ the $\arg\!\min$ of the distilled DPO objective (eq. 7) is $\min(1-\hat{\pi}_{\theta}(y_{i}^{w}),\hat{\pi}_{\theta}(y_{i}^{\ell})$, with $\log\hat{\pi}_{\theta}(y_{i}^{\ell})=\frac{1}{\beta}(r_{t}(x,y_{i}^{\ell})-r_{t}(x,y_{i}^{w}))+\log\hat{\pi}_{\theta}(y_{i}^{w})+\log\frac{\pi_{\mathrm{ref}}(y_{i}^{\ell})}{\pi_{\mathrm{ref}}(y_{i}^{w})}$.
@@
Why DPO can drive $\pi(y^{w})$ to zero.
In §2.3 we pointed out a peculiarity of the DPO global optima: in certain cases, it can include policies where $\pi(y^{w})$ may be nearly 00 for all $y^{w}$ in the training set. This undesirable behavior has also been observed in practice [20; 22; 30]. For intuition on why this may happen, consider the simplified case where the policy is a bag-of-words model, $\pi_{\theta}(y)\propto\exp\left(c(y)\cdot\theta\right)$ for $c(y)$ representing a vector of counts in $y$ and $\theta_{i}$ representing the unnormalized log-probability of token $i$. Then we can formally show that DPO optimization monotonically decreases an upper bound on the probability of the preferred completion, $\tilde{\pi}_{\theta^{(t-1)}}(y^{w})\geq\tilde{\pi}_{\theta^{(t)}}(y^{w})\geq\pi_{\theta^{(t)}}(y^{w}).$
@@
Let $y^{w},y^{\ell}\in\mathcal{V}^{n}$ be preferred vs. dispreferred outputs of length $n$, respectively, with $\pi_{\mathrm{ref}}(y^{w}),\pi_{\mathrm{ref}}(y^{\ell})>0$ and corresponding count vectors $c(y^{w}),c(y^{\ell}).$ Let $\log\pi_{\theta}(y)=c(y)\cdot\theta-nZ(\theta)$ for $Z(\theta)=\log\sum^{\mathcal{V}}_{i}e^{\theta_{i}}$, with upper bound $\log\tilde{\pi}_{\theta}(y)=c(y)\cdot\theta-n\max_{j}\theta_{j}$. Let $\theta^{(t)}$ represent the parameters of $\pi$ after $t$ steps of gradient descent on $\mathcal{L}_{\mathrm{dpo}}(\{y^{\ell},y^{w},x\})$, with $\theta^{(0)}=0$. Then $\pi_{\theta^{(t)}}(y^{w})\leq\tilde{\pi}_{\theta^{(t)}}(y^{w})\leq\tilde{\pi}_{\theta^{(t-1)}}(y^{w})$ for all $t$.
Where does the probability mass go?
If $\pi_{\theta^{(t)}}(y^{w})$ decreases in $t$, what other strings become more probable? In the following proposition, we show that under the bag-of-words model, DPO optimization moves probability mass away from $y^{w}$ to sequences that contain only the tokens that maximize the difference between $y^{w}$ and $y^{\ell}$. This is a concrete example of the type of undesirable optima described in §2.3, now shown here to be realizable.
Proposition 6.
Let $y^{w}$ and $y^{\ell}$ be preferred / dispreferred outputs of length $n$. Let $\Delta=c(y^{w})-c(y^{\ell})$ be the difference in unigram counts. Let $\hat{y}=[i,i,\ldots,i]$, for $i\in\text{arg}\max\Delta,$ with $\|c(\hat{y})\|_{1}=n$. Then $\pi_{\theta^{(t)}}(y^{w})-\pi_{\theta^{(t)}}(\hat{y})=\tau(t)k$ for some $k\leq 0$ and some non-decreasing $\tau:\mathbb{Z}_{+}\to\mathbb{R}_{+}$.
We have $k=0$ when $c(y^{w})=c(\hat{y})$, and $k\ll 0$ when $\|c(y^{w})\|_{2}\ll\|c(\hat{y})\|_{2}=n$ (dense $c(y^{w})$) and $\|\Delta\|_{2}=\|\Delta\|_{\infty}$ (sparse $\Delta$). This implies that when $y^{w}$ and $y^{\ell}$ are similar, $\pi_{\theta}(y^{w})$ will degrade more rapidly. Early stopping will therefore tradeoff between reaching the degenerate solution on such cases, and underfitting other cases in which $y^{w}$ and $y^{\ell}$ are more distinct.
6 Related work
Recent work in offline alignment has focused on DPO [23] as a simpler alternative for aligning language models from preference data. Subsequent work has identified issues with DPO, including weak regularization [3] and a tendency to decrease the probability of winning generations during training [20]. Other methods have explored various avenues for improvement. These include analyzing the impact of noise on DPO alignment [11], proposing to update the reference policy during training [12], and suggesting a variant of IPO with a per-context margin [1]. Additional research has focused on token-level alignment methods [38; 22] and on developing a unified view of various offline alignment methods [31]. This work builds upon several these findings, and provides further analysis, as well as a solution based on pessimism and reward distillation.
While offline alignment methods are popular, recent evidence suggests that online alignment methods such as RLHF [6; 29], may lead to more favorable outcomes [13; 30; 8; 34]. Notably, Zhu et al. [41] proposed iterative data smoothing, which uses a trained model to softly label data during RLHF. Whether online or offline, however, policies are still succeptible to overfitting to certain degenerate phenomena. To this end, reward ensembles have been widely investigated recently as a mechanism for tackling reward hacking in RLHF [9; 7; 39; 25], and in the context of multi-objective optimization [19; 24]. We use an ensemble of rewards to represent the uncertainty with respect to reward models that are suitable given preference data. Moskovitz et al. [19] focus on “composite” rewards, with the goal of achieving high task reward while ensuring that every individual component is above some threshold—also by applying a Lagrangian relaxation. In this work, we also consider multiple reward models, but we only focus on cases where there is no known, obvious reward decomposition.
Finally, the question of using a small amount of offline data to learn high-quality policies, instead of online access to reward feedback, has been widely studied in the offline reinforcement learning (RL) literature. The predominant approach here is to use pessimism, that is, to learn a policy with the highest reward under all plausible environment models consistent with the data, with an extensive theoretical [18; 37; 33] and empirical [16; 5; 36] body of supporting work. The key insight in this literature is that without pessimism, the RL algorithm learns undesirable behaviors which are not explicitly ruled out in the training data, and pessimism provides a robust way of preventing such undesirable extrapolations, while still preserving generalization within the support of the data.
@@ 7 Conclusion
LM alignment is crucial for deploying safe and helpful assistants, but is difficult due to lack of access to perfect preference oracles. We presented a thorough theoretical analysis of some of the degeneracies that DPO is susceptible to when learning from sampled human preference data. Furthermore, our findings suggest that explicit reward modeling remains a powerful vehicle for introducing regularization into post-training. By distilling the reward assigned by a single, explicit reward model—or a family of explicit reward models—directly into the implicit reward maximized by our policies using offline data, we demonstrated that we can achieve improved robustness to variations in preference dataset quality, while maintaining the simplicity of the DPO framework.
Limitations. The empirical results in the paper are based on one dataset and form of distribution shift. For deeper understanding of pessimism and ensembling, additional settings should be explored. The theoretical aspects of the paper are sometimes based on restrictive assumptions and simplifications. Nonetheless, they provide potential explanations for phenomena observed in real-world settings.
Broader impact. We introduce new ideas to the active field of research on preference-based post-training, which we hope will help facilitate the alignment of large models, and improve understanding of current approaches—ultimately supporting the development of capable and reliable AI systems.
Acknowledgements. We thank Alexander D’Amour and Chris Dyer for helpful comments and feedback on the manuscript. This research also benefited from discussions with Victor Veitch, Mandar Joshi, Kenton Lee, Kristina Toutanova, David Gaddy, Dheeru Dua, Yuan Zhang, Tianze Shi, and Anastasios Angelopoulos.
@@ A.5 Proof of Proposition 2
A.6 Proof of Proposition 3
@@
A.7 Proof of Proposition 4
@@
This follows directly from differentiating eq. 7 with respect to $\pi_{\theta}(y_{2}).$
A.8 Proof of Proposition 5
Proof.
$\frac{\partial\mathcal{L}_{\beta}(\theta)}{\partial\theta}=$ $-(1-\sigma(\beta(\Delta\cdot\theta)+\beta\log\rho))\beta\Delta=-\Pr(y^{\ell}\succ y^{w};\theta)\beta\Delta\prec 0.$ (37)
Let $\delta_{t}=\beta\Pr(y^{\ell}\succ y^{w};\theta^{(t)}).$ Then,
A.9 Proof of Proposition 6
Proof.
$\theta^{(t)}=$ $\theta^{(t-1)}+\eta\beta\Pr(y^{\ell}\succ y^{w};\theta^{(t-1)})\Delta=\left(\sum^{t}_{t^{\prime}=1}\eta\beta\Pr(y^{\ell}\succ y^{w};\theta^{(t^{\prime})})\right)\Delta=\tau(t)\Delta.$ (43)
Plugging these parameters into the likelihoods,
$=\tau(t)(c(y^{w})\cdot\Delta-n\max\Delta)=\tau(t)k,$ (46)
Appendix B Transitive closure
Both p-DPO and IPO target a constant ratio for $\log\psi_{w}/\psi_{l}$. However, the loss surfaces are different. To see this, we consider a simplified setting with three possible outputs, $y_{1}$, $y_{2}$, $y_{3}$. We observe either $\mathcal{D}=\{(y_{1}\prec y_{2}),(y_{2}\prec y_{3})\}$ or $\overline{\mathcal{D}}=\mathcal{D}\cup\{(y_{1}\prec y_{3})\}$. If we treat this problem as a multi-arm bandit, the goal is to assign a weight to each arm, which we denote $\psi_{i}=\log\pi_{\theta}(y_{i}|x)+Z_{x},$ with $Z_{x}$ an underdetermined log-partition function.
Proposition 7.
@@ Appendix B Transitive closure
Intuitively, the observation of ${(y_{1}\prec y_{3})}$ should increase confidence that $y_{3}$ is superior to $y_{1},$ but in IPO it has the opposite effect, drawing their scores closer together. While pessimistic DPO also has a target ratio between each preference pair, its loss surface is different: in particular, it does not increase quadratically as we move away from the target. We find empirically that pessimistic DPO is robust to the transitive closure of preference annotations in the multi-arm bandit setting, as shown in Figure B.1. As discussed above, DPO will set $\psi_{1}\to-\infty$ because $y_{1}$ is never preferred.
In our experiments we solve the p-DPO and IPO objectives for both $\mathcal{D}=\{(y_{1},y_{2}),(y_{2},y_{3})\}$ and $\mathcal{\overline{D}}=\mathcal{D}\cup\{(y_{1},y_{3})\}$, solving with respect to $\{\pi_{\theta}(y_{i})\}.$ IPO is solved analytically as a quadratic program; for p-DPO we used projected gradient descent. We consider $\beta\in(1,3,10,30)$ and $\alpha\in(5,10,20,50,100,1000)$. As shown in Figure B.1, there are significant differences in the IPO solutions with and without transitive closure, while for p-DPO these differences are imperceptible.
@@
Appendix C Distribution over reward models for e-DPO
@@
Figure C.1: We show for every length bias, $\rho$, the distribution over reward models that best match the final policy trained by e-DPO across all training examples. We observe that the e-DPO policy matches different reward models across examples. Moreover, when the policy is trained with data biased towards preferring short responses, the reward model that was trained on longer responses is often preferred and vice versa.
Appendix D Hyperparameters
Validation set performance across the range of hyperparameter settings is shown in Figure D.1. In pilot studies we found that these results were relatively robust to variation in the random seed, but did not conduct extensive investigation of this effect across all methods and hyperparameters due to cost.
Figure D.1: Validation set results across hyperparameters for each method. For all methods, different values of $\rho$ induce different optimal hyperparameters $\beta$ and $\tau^{-1}$.
Appendix E Compute resources

### v2

@@ 7.1 Optima
The optimum in Proposition 5 corresponds to $\log\psi_{w}/\psi_{\ell}={\beta}^{-1}\log(\alpha-1)$. Recall that IPO Azar et al. (2024) seeks to assign a constant value to this ratio by minimizing $(\log\frac{\psi_{w}}{\psi_{\ell}}-\tau^{-1})^{2}$; the (unconstrained) optima are identical for $\tau^{-1}:={\beta}^{-1}\log(\alpha-1)$, but the loss surfaces are different (see further analysis of this in §7.2). DPO sets $\pi_{\theta}(y^{\ell}_{i})\to 0$, as shown in Corollary 1; this is due not only to competition from $\pi_{\theta}(y^{w}_{i})$ but from DPO penalizing positive probability on $y^{\ell}_{i}$. Analysis of the distilled loss gives a similar result:
@@
For any fixed $\hat{\pi}_{\theta}(y_{i}^{w})$ and $\beta>0,$ the $\arg\!\min$ of the distilled DPO objective in (7) is $\min(1-\hat{\pi}_{\theta}(y_{i}^{w}),\hat{\pi}_{\theta}(y_{i}^{\ell}))$, with $\log\hat{\pi}_{\theta}(y_{i}^{\ell})=\frac{1}{\beta}(r_{t}(x,y_{i}^{\ell})-r_{t}(x,y_{i}^{w}))+\log\hat{\pi}_{\theta}(y_{i}^{w})+\log\frac{\pi_{\mathrm{ref}}(y_{i}^{\ell})}{\pi_{\mathrm{ref}}(y_{i}^{w})}$.
@@
7.2 Transitive closure: p-DPO vs IPO
As pointed out in §7.1, both p-DPO and IPO target a constant ratio for $\log\psi_{w}/\psi_{l}$. Despite the similar form of optima, however, the loss surfaces of the two objectives differ in notable ways. To see this, we consider a simplified setting with three possible outputs, $y_{1}$, $y_{2}$, $y_{3}$. We observe either $\mathcal{D}=\{(y_{1}\prec y_{2}),(y_{2}\prec y_{3})\}$ or $\overline{\mathcal{D}}=\mathcal{D}\cup\{(y_{1}\prec y_{3})\}$. If we treat this problem as a multi-arm bandit, the goal is to assign a weight to each arm, which we denote $\psi_{i}=\log\pi_{\theta}(y_{i}\mid x)+Z_{x},$ with $Z_{x}$ an underdetermined log-partition function.
@@ 7.2 Transitive closure: p-DPO vs IPO
Intuitively, the observation of ${y_{1}\prec y_{3}}$ should increase confidence that $y_{3}$ is superior to $y_{1},$ but in IPO it has the opposite effect, drawing their scores closer together. While pessimistic DPO also has a target ratio between each preference pair, its loss surface is different: in particular, it does not increase quadratically as we move away from the target. We find empirically that pessimistic DPO is robust to the transitive closure of preference annotations in the multi-arm bandit setting, as shown in Figure B.2. As discussed above, DPO will set $\psi_{1}\to-\infty$ because $y_{1}$ is never preferred. Specifically, in our experiments we solve the p-DPO and IPO objectives for both $\mathcal{D}=\{(y_{1},y_{2}),(y_{2},y_{3})\}$ and $\mathcal{\overline{D}}=\mathcal{D}\cup\{(y_{1},y_{3})\}$, solving with respect to $\{\pi_{\theta}(y_{i})\}.$ IPO is solved analytically as a quadratic program; for p-DPO we used projected gradient descent. We consider $\beta\in(1,3,10,30)$ and $\alpha\in(5,10,20,50,100,1000)$. As shown in Figure B.2, there are significant differences in the IPO solutions with and without transitive closure, while for p-DPO these differences are imperceptible.
Takeaways: Theoretical analysis Both pessimistic DPO and reward model distillation avoid the degenerate optima of DPO that were analyzed in §4—primarily through adding additional sources of regularization to the objective. In addition to these more stable optima, we also show how the loss surface of p-DPO can lead to more favorable outcomes with respect to modeling challenging transitive preferences vs. IPO.
@@ 8 Conclusion
LM alignment is crucial for deploying safe and helpful assistants, but is difficult due to lack of access to perfect preference oracles. We presented a thorough theoretical analysis of some of the degeneracies that DPO is susceptible to when learning from sampled human preference data. Furthermore, our findings suggest that explicit reward modeling remains a powerful vehicle for introducing regularization into post-training. By distilling the reward assigned by a single, explicit reward model—or a family of explicit reward models—directly into the implicit reward maximized by our policies using offline data, we demonstrated that we can achieve improved robustness to variations in preference dataset quality, while maintaining the simplicity of offline alignment frameworks. Finally, reward model distillation also results in modest but consistent improvements in performance even on unbiased settings, making it an overall compelling algorithmic modification to offline training.
Acknowledgements
We thank the anonympous reviewers, Alexander D’Amour, and Chris Dyer for helpful comments and feedback on the manuscript. This research also benefited from discussions with Victor Veitch, Mandar Joshi, Kenton Lee, Kristina Toutanova, David Gaddy, Dheeru Dua, Yuan Zhang, Tianze Shi, and Anastasios Angelopoulos.
@@ A.4 Proof of Theorem 2
Theorem (Theorem 2 restated).
Define the constrained minimizer
$\pi_{\theta^{*}}(y\mid x)\in\underset{\pi_{\theta}\in\mathcal{P}_{\beta}(\mathcal{S})}{\arg\!\min}~{}\beta\mathbb{E}_{\mu(x)}\mathbb{D}_{\mathrm{KL}}(\pi_{\mathrm{ref}}(\cdot\mid x)\|\pi_{\theta}(\cdot\mid x)),$
@@ A.5 Proof of Proposition 2
Proposition (Proposition 2 restated).
Let $y^{w},y^{\ell}\in\mathcal{V}^{n}$ be preferred versus dispreferred outputs of length $n$, respectively, with $\pi_{\mathrm{ref}}(y^{w}),\pi_{\mathrm{ref}}(y^{\ell})>0$ and corresponding count vectors $c(y^{w}),c(y^{\ell}).$ Let $\log\pi_{\theta}(y)=c(y)\cdot\theta-nZ(\theta)$ for $Z(\theta)=\log\sum^{\mathcal{V}}_{i}e^{\theta_{i}}$, with upper bound $\log\tilde{\pi}_{\theta}(y)=c(y)\cdot\theta-n\max_{j}\theta_{j}$. Let $\theta^{(t)}$ represent the parameters of $\pi$ after $t$ steps of gradient descent on $\mathcal{L}_{\mathrm{dpo}}(\{y^{\ell},y^{w},x\})$, with $\theta^{(0)}=0$. Then, we have that
$\pi_{\theta^{(t)}}(y^{w})\leq\tilde{\pi}_{\theta^{(t)}}(y^{w})\leq\tilde{\pi}_{\theta^{(t-1)}}(y^{w})\quad\text{for all $t$,}$
with strict inequality when $||c(y^{w})-c(y^{\ell})||_{0}>1$.
Proof.
$\frac{\partial\mathcal{L}_{\beta}(\theta)}{\partial\theta}=$ $-(1-\sigma(\beta(\Delta\cdot\theta)+\beta\log\rho))\beta\Delta=-p(y^{\ell}\succ y^{w};\theta)\beta\Delta\prec 0.$ (30)
Let $\delta_{t}=\beta p(y^{\ell}\succ y^{w};\theta^{(t)}).$ Then,
A.6 Proof of Proposition 3
Proposition (Proposition 3 restated).
Let $y^{w}$ and $y^{\ell}$ be preferred versus dispreferred outputs of length $n$. Let $\Delta=c(y^{w})-c(y^{\ell})$ be the difference in unigram counts. Let $\hat{y}=[i,i,\ldots,i]$, for $i\in\text{arg}\max\Delta,$ with $\|c(\hat{y})\|_{1}=n$. Then $\pi_{\theta^{(t)}}(y^{w})-\pi_{\theta^{(t)}}(\hat{y})=\tau(t)k$ for some $k\leq 0$ and some non-decreasing $\tau:\mathbb{Z}_{+}\to\mathbb{R}_{+}$.
Proof.
$\theta^{(t)}=$ $\theta^{(t-1)}+\eta\beta p(y^{\ell}\succ y^{w};\theta^{(t-1)})\Delta=\left(\sum^{t}_{t^{\prime}=1}\eta\beta p(y^{\ell}\succ y^{w};\theta^{(t^{\prime})})\right)\Delta=\tau(t)\Delta.$ (36)
Plugging these parameters into the likelihoods,
$=\tau(t)(c(y^{w})\cdot\Delta-n\max\Delta)=\tau(t)k,$ (39)
A.7 Proof of Proposition 4
Proposition (Proposition 4 restated).
Assume the same conditions as Theorem 1. Then for any $0<\gamma<\infty$, there exists a $\lambda\geq 0$ such that $\pi_{\theta^{*}}(y\mid x)\in\arg\!\min_{\pi_{\theta}}\mathcal{L}_{\mathrm{pdistill}}(\mathcal{S},\pi_{\theta};\rho)$, where $\pi_{\theta^{*}}$ is a minimizer over all possible policies of the objective (9), for the effective reward model set
$\mathcal{S}_{\gamma}=\bigcup_{r_{\mathrm{tgt}}^{i}\in\mathcal{S}}\Big{\{}\tilde{r}\colon\mathbb{E}_{\rho(x,y_{1},y_{2})}\left[(r_{\mathrm{tgt}}^{i}(x,y_{1})-r_{\mathrm{tgt}}^{i}(x,y_{2})-\tilde{r}(x,y_{1})+\tilde{r}(x,y_{2}))^{2}\right]\leq\lambda\Big{\}}.$
@@ A.7 Proof of Proposition 4
A.8 Proof of Proposition 5
Proposition (Proposition 5 restated).
@@ A.8 Proof of Proposition 5
A.9 Proof of Proposition 6
Proposition (Proposition 6 restated).
For any fixed $\hat{\pi}_{\theta}(y_{i}^{w})$ and $\beta>0,$ the $\arg\!\min$ of the distilled DPO objective in (7) is $\min(1-\hat{\pi}_{\theta}(y_{i}^{w}),\hat{\pi}_{\theta}(y_{i}^{\ell}))$, with $\log\hat{\pi}_{\theta}(y_{i}^{\ell})=\frac{1}{\beta}(r_{t}(x,y_{i}^{\ell})-r_{t}(x,y_{i}^{w}))+\log\hat{\pi}_{\theta}(y_{i}^{w})+\log\frac{\pi_{\mathrm{ref}}(y_{i}^{\ell})}{\pi_{\mathrm{ref}}(y_{i}^{w})}$.
@@ A.9 Proof of Proposition 6
This follows directly from differentiating (7) with respect to $\pi_{\theta}(y_{2}).$
A.10 Proof of Proposition 7
Proposition (Proposition 7 restated).
@@ A.10 Proof of Proposition 7
Appendix B Additional results
B.1 The effect of distillation and pessimism on likelihood collapse
Figure B.1 tests the effect of pessimism in both DPO (DPO vs p-DPO) and distilled DPO (d-DPO vs dp-DPO) on likelihood collapse in the preferred vs. dispreferred generations in the offline data. Note that a subtle point arises from the distinction between regularizing to the reference policy output distribution versus regularizing to the preference data distribution, which are only (asymptotically) identical if preferences are annotated on a sample from the reference policy (in our initial experiments we had found slightly better results overall when regularizing to the reference policy output distribution). Figure B.1 reports results with respect to both distributions, showing that pessimism (a) mitigates the decrease in probability of preferred and dispreferred preference annotations, despite this data not being used in the regularizers (left-most subplots), and (b) mitigates the increase in KL divergence with respect to the reference distribution (third subplot from left), as expected due to the additional regularization term. As argued in §4 (see also Azar et al. [2024]), the $\beta$ hyperparameter of DPO does not effectively regularize this KL distribution because the implicit DPO reward model assigns infinite-magnitude rewards.
Figure B.1: Pessimism mitigates likelihood collapse. In this case, we penalize the sample KL-divergence rather than the data KL-divergence, because the samples are not drawn from the preference data distribution. Thus the effect of regularization is primarily on kl_loss_ref, which quantifies this divergence, rather than on the likelihood of the preference data itself, (dis)preferred_log_likelihood.
B.2 Transitive closure
Figure B.2 shows the results of our multi-arm bandit experiments with p-DPO vs. IPO losses, as described in §7.2. In this synthetic setup, we solve the p-DPO and IPO objectives for both $\mathcal{D}=\{(y_{1},y_{2}),(y_{2},y_{3})\}$ and $\mathcal{\overline{D}}=\mathcal{D}\cup\{(y_{1},y_{3})\}$, solving with respect to $\{\pi_{\theta}(y_{i})\}.$
@@ B.2 Transitive closure
B.3 Distribution over reward models for e-DPO
@@
Figure B.3: For all training examples, we track which reward model during training best matches the implicit reward of the current e-DPO policy and plot the distribution over reward models, for every length bias, $\rho$. We observe that the e-DPO policy matches different reward models across examples during training. Moreover, when the policy is trained with data biased towards preferring short responses, the reward model that was trained on longer responses is by and large preferred and vice versa.
B.4 Hyperparameters
Validation set performance across the range of hyperparameter settings is shown in Figure B.4. Figure B.5 also explores the impact of $\gamma$ (which is minimal). In pilot studies we found that these results were relatively robust to variation in the random seed, but did not conduct extensive investigation of this effect across all methods and hyperparameters due to cost.
Figure B.4: Validation set results across hyperparameters for each method. For all methods, different values of $\rho$ induce different optimal hyperparameters $\beta$ and $\tau^{-1}$.
Figure B.5: Comparing e-dpo with and without annealing of $\gamma$ on the development set for all length biases using the best value of $\beta$ per bias. Annealing $\gamma$ has minimal effect on performance and is not necessary for the success of e-dpo.
Appendix C Compute resources